A sample of 4.6434 g of benzoic acid, C_7H_6O_2, i
Answer #1

Mw benzoic acid = 122.12

wt benzoic acid = 4.6434g

Now, moles of benzoic acid = 4.6434g / 122.12 = 0.03801

Depression in freezing point = Kf X moles of solute (benzoic acid) X 1000 / wt of solvent (camphor)

Depression in freezing point = 40 X 0.3801 X 1000 / 50.0033

Depression in freezing point = 30.41 oC

From now, Depression in freezing point = Tsolvent - Tsolution

or, Tsolution = Tsolvent - Depression in freezing point

    Tsolution = 179.5 - 30.41 = 149.09 oC